OpenStudy (clamin):

You want to use the hot tub at the hotel you're staying at, but it's not hot enough. The hotel manager tells you they they will turn the heat up and the temperature will increase by 10% each hour.The initial temperature of the hot tub is 68° If the temperature increases by 10% eaxh hour, what is the common ratio?

OpenStudy (welshfella):

convert 10% to decimal form. That is your common ratio.

OpenStudy (clamin):

So its 0.1?? @welshfella

OpenStudy (welshfella):

yes

OpenStudy (clamin):

So if i solve it with explicit formula what is my a1?

OpenStudy (clamin):

@welshfella

OpenStudy (welshfella):

- I apologies 0.10 is incorrect it should be 1.1 so the first few terms of the sequence of temperatures is 68 , 68*1.1 , 68 * 1.1^2 = 68 , 74.8, 82.3

OpenStudy (welshfella):

if a1 means the first term its 68 second term is 74.8 degrees

OpenStudy (clamin):

Why is it 1.1?? @welshfella

OpenStudy (welshfella):

well 0.10 is the fractional INCREASE the next temperatuer in sequence is 68 + 89* 0.1 = 68*1.1 this is a geometric sequence with common ratio 1.1

OpenStudy (welshfella):

* typo 68+68*0.1

OpenStudy (clamin):

Thanks can you help also with my another question its in math sectiom

OpenStudy (welshfella):

i'll take a look